OpenStudy (anonymous):

How many units apart are the points (4, -3) and (-4, 3)?

OpenStudy (anonymous):

to find the distance (or units) apart between two points \[(x_1,y_1) and (x_2,y_2)\] \[D=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

D is the Distance apart between the two points...

OpenStudy (anonymous):

what do you get if you use the formula??

OpenStudy (anonymous):

um.. so far 0+-8. Am I wrong?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

T-T I suck at this...

OpenStudy (anonymous):

i think the mistake is from..... x(2)-x(1) so it is supposed to be (4-(-4))^2

OpenStudy (anonymous):

Oh I think I understand...

OpenStudy (anonymous):

-16?

OpenStudy (anonymous):

no..its supposed to be 64 for (4-(-4))^2

OpenStudy (anonymous):

do you understand why??

OpenStudy (anonymous):

4-(-4)=4+4=8 and therefore 8^2=64

OpenStudy (anonymous):

oh. Thank you very much lol I feel dumb

OpenStudy (anonymous):

so what do you get when you find the distance??

OpenStudy (anonymous):

Let d be the distance between the points.Then \[d=\sqrt{\left( -4-4 \right)^{2}+\left( 3+3 \right)^{2}}\] \[d=\sqrt{64+36}=\sqrt{100}=10\]